satellite73 (satellite73):

anyone feel like computing an integral using residues? \[\int_0^{2\pi}\frac{d\theta}{a+\sin(\theta)},a\in \mathbb R, a>1\]gimmick is to write \[z=e^{i\theta}\] and then use residues.

OpenStudy (anonymous):

guess not. i don't feel like it either, but i guess i am stuck with it

OpenStudy (anonymous):

ok it wasn't that bad after all

OpenStudy (anonymous):

think the answer is (2*pie/(a^2 -1)^0.5) .. I used the following algorithm .. algorithm is sin(x) = (e^(ix) - e^(-ix))/2 hence we have to find poles for (2/((z^2 -1) +2*i*a*z)) poles would be z = i*(-a +- (a^2-1)^0.5) only z =p1 = i*(-a + (a^2 -1)^(0.5)) would lie in the region z = e^(ix) .. hence find the limit z -> p1 for (z - p1)*f(z) = (1/i*(a^2-1)^0.5) now simply multiply by 2*pie*i to get 2*pie/(a^2-1)^0.5